NATIONAL RADIO ASTRONOMY OBSERVATORY
Green Bank, WV
October 12, 1999
TO: R. Fisher, J. Ford, F. Ghigo, R. Lacasse, C. Niday, T. Weadon
FROM: R. Lacasse
SUBJ.: Minutes of Interference Protection Group 10/12/99 Meeting
In attendance: Fisher, Lacasse, Niday
Minutes of Previous Meeting
Call Outs for Interference
Action Item Review
1. Automatic Paging
The OVLBI group has requested that an automatic paging system be used
to contact personnel in their group when errors occur in their system.
The paging system would use the site 43 MHz, 60-W transmitter. It
would be capable of penetrating buildings. Bignell or Pedtke was
supposed to contact Niday about testing such a system. Nothing has
happened so far. Possibly the concept is on hold for lack of someone
to work on it.
2. Turntable for the
Egan has located a unused turntable near the water tower. It
is fairly massive, but might be suitable for use in the chamber if it were
modified. Niday inspected it and reported that it does not seem practical.
It would be a lot of work to modify and would have to be a permanent fixture
in the chamber because it is so heavy. Fisher will report back to
3. Allocation of Remaining
IPG Funds for 1999
Approximately $1500 was spent in Monitor Station parts. Most
of this year's money is now spent, although a few hundred dollars could
be found for shielding/filtering projects if necessary.
from the 140-ft. for the SOHO Mission
The possibility of using the 140-ft. telescope as a ground station
for the SOHO mission was brought up recently at a scientific staff meeting.
During the past month quite a bit of correspondence by e-mail has taken
place. As a result, some recommendations from the IPG were formulated.
Fisher believes that these recommendations were incorporated in the proposal
sent to NASA.
5. Site radio system
The antennas are scheduled to be installed on the water tower tomorrow.
6. Microwave oven in shielded
test box (John)
The NRAO panel setting crew is very interested in using the system.
The NRAO crew installing cryo lines on the GBT feed arm is presently
using the portable units for line of sight communications and is
happy with them.
7. RFI Considerations for
the New Tour Center
8. Shielded Room Entrances
9. Monitor Station Developments
As reported above, the last of this year's funds have been spent on
this project. Labor is scarce right now, so progress on construction
is very slow.
10. Progress on Suppressing RFI
from Equipment Measured in Anechoic Chamber
Schiebel is continuing to test components for the feed defroster.
No progress on
other items on action item list.
Apparently there has been no progress in the laser group in this regard.
Lacasse reported that progress has been made in getting the Spectral
Processor back on the air in the GBT equipment room. All copper cabling
has been completed. Fiber cables for SIG/REF and IF from the anechoic
chamber to the spectral processor are on Childers' list of cables to run.
There is a problem with the 1 PPS going to the rack: the trailing edge
rings quite a lot. This is true for all copies of the 1 PPS in the room.
The processor does not start scans properly, and this may be related to
the 1 PPS. Fixing this is a fairly high priority.
An article from the September 1999 RFI Design journal, entitled Basic
Steps to Successful EMC Design, was distributed to those in attendance.
It was decided to make copies for all engineers an technicians.
Fisher reported that he had recently been to Arecibo for an observing
run. He reported that a birdie from the LO system was present at
1400 MHz. This should be good motivation for people in Green Bank
to try to avoid similar problems. He also took advantage of time
there to check on what Arecibo is doing for RFI mitigation. They
have one full time person who runs a monitoring station. The station
covers 800 to 5000 MHz, and perhaps a few spot frequencies below 800 MHz.
Much of the effort is directed towards monitoring the spectrum. Their
biggest problem is sensitivity, i.e., the monitoring station cannot see
everything that the astronomy system does.
Next Meeting: October 26, 0800.